40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 13720     Entry time: Fri Mar 30 03:23:50 2018
Author: Koji 
Type: Update 
Category: General 
Subject: aLIGO EOM work 

I have been working on the aux beat setup on the PSL table between 9PM-3AM.

This work involved:

- Turning off the main marconi
- Turning off the freq generation unit (incl IMC modulation)
- Closing the PSL shutter

After the work, these were reverted and the IMC and both arms have been locked.

ELOG V3.1.3-